# Hybrid Architecture for Real-Time Video Anomaly Detection: Integrating Spatial and Temporal Analysis

### v1

@@ 5 Conclusion
5.1 Limitations and Future Work
Our study has highlighted several challenges and improvement perspectives for our anomaly detection system. Firstly, we observed that not all anomalies are necessarily linked to identifiable key objects, particularly in the case of natural disasters. Moreover, the presence of key objects does not always signify danger, as illustrated by the example of armed military personnel in airports. To overcome these limitations, we are considering directly transmitting the information collected by YOLO and VGG+GRU to our MLP, allowing it to automatically learn the conditions for anomaly detection. We also noticed confusion between certain anomaly classes, suggesting the potential benefit of exploring a binary model to verify this hypothesis. Another promising avenue would be to combine several different processing methods, particularly by only showing detected objects and representing people by their skeletons, which could improve the reliability of the serial model, albeit at the expense of detection speed. Finally, to enhance the robustness and versatility of our system, it would be wise to enrich our dataset with new anomaly classes and compare our architecture to other existing models in the field.

### v3

@@ 6 Conclusion and Future works
Our study has highlighted several challenges and improvement perspectives for our anomaly detection system. Firstly, we observed that not all anomalies are necessarily linked to identifiable key objects, particularly in the case of natural disasters. Moreover, the presence of key objects does not always signify danger, as illustrated by the example of armed military personnel in airports. To overcome these limitations, we are considering directly transmitting the information collected by YOLO and VGG+GRU to our Multi-Layer Perceptron, allowing it to automatically learn the conditions for anomaly detection. We also noticed confusion between certain anomaly classes, suggesting the potential benefit of exploring a binary model to verify this hypothesis.
Another promising approach would be to combine several different processing methods. For example, by associating a mask with the detected objects and representing people by their skeletons, which could improve the reliability of the sequential model, at the cost of reduced detection speed. Finally, to enhance the robustness and versatility of our system, it would be beneficial to enrich our dataset with new anomaly classes and examine the execution speed of our architecture compared to other existing models in the field.